1BF the work you have completed so far, it does not indicate come 10 Problem 4-22 Interest-on-Interest (LG4-3) 2.5 points Con what is part two
Answer #1

a.

FV = PV(1+r)n

= 4,800 * (1+0.10)10

FV = $12,449.95

b.

Interest Earned = Amount - Principal

= 12,449.95 - 4,800

Interest Earned = $7,649.95